A 75-kg man pushes the same cart up a 10-m ramp elevated an angle of 15O. Find the work done.
1
Expert's answer
2021-02-24T12:49:28-0500


I assume, he pushes the cart with constant velocity up along the ramp.

Then the force "\mathbf{F}" applied to the cart by the men is equal to the friction force "\mathbf{F}_{fr}" along the ramp and the projection of the gravitational force "m\mathbf{g}" (see figure). Thus, obtain:


"F = F_{fr} + mg\sin\alpha"

where "m" is the mass of the cart together with the man (assuming the task is to find all work done; if the task is to find the work done with respect to the cart, then "m" will be the mass of the cart only), and "\alpha = 15\degree" is the angle of inclination. By definition, the frictional force is:


"F_{fr} = \mu N"

where "\mu" is the coefficient of friction, and "N = mg\cos \alpha" is the normal force (see figure). Thus, obtain:


"F = \mu mg\cos \alpha + mg\sin \alpha = mg(\mu \cos \alpha +\sin \alpha)"

By difinition, the work is:


"A = Fs"

where "s = 10m" is the length of the ramp. Finally, obtain:


"A = smg(\mu \cos \alpha +\sin \alpha)"

As one can see, there are several quantities, that are omitted in this prombem (namely, "m,\mu"). Thus, it is not possible to obtain a numerical answer. But the final formula was derived.


Answer. "A = smg(\mu \cos \alpha +\sin \alpha)".